if a olympic swimming pool is twice as long as it is wide and ther perimeter is 150 meters then what are the lenght and width?
Answer by mananth(16946)

Let width be x
Length = 2x
..
Perimeter = 2*(L+B)
150= 2(x+2x)
150=6x
x=150/6
x=25 meters width
Length = 2x = 50 meters